OpenStudy (anonymous):

You spin the pointer of this spinner. Find the probability that the pointer will stop on a number greater than 6 The spinner ranges from 1 till 8

OpenStudy (anonymous):

2/8 =1/4 =.25

OpenStudy (anonymous):

NADIAHATESCESAR I THINK AN94 IS RIGHT

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

thanks